OpenStudy (anonymous):

What is the best definition for: 1) Capacitance 2) Inductance

OpenStudy (anonymous):

Capacitance of a conductor is the ability to store charges and is calculated as the ratio of charge to the potential the meaning is, how well can a conductor store charges without raising its potential much why is that important? think of charges as water, and the height of the water as the potential.. if a container can store water with less height, it definitely means it has a high capacity right? so if a conductor can store charges at low potential then it has high capacity... why we need high capacity?? what happens if conductors store charges at high potential?? then just like a container, ultimately the water spills out, here also the charges will leak out if potential goes too high.. hope that tells you about capacitance :P.. if you like it then ll tell you about inductance as well

OpenStudy (chenping):

Mathematically, C=Q/V and so theoritically it's as he said above. I will talk about Inductance then ... :P It's name originates from Inductor, a circuit element, typically a conducting coil, in which electromotive force is generated by electromagnetic induction. So, inductance is basically the property of an electric circuit by which an electromotive force is induced in it (both the conductor and inductor) as the result of a changing magnetic flux. \displaystyle v= L\frac{di}{dt} where v denotes the voltage in volts and i the current in amperes.

OpenStudy (chenping):

sorry, equation error. It's v=L(dI/dt)